OpenStudy (anonymous):

96 / 4^2 * 3^2 + 2^3

OpenStudy (anonymous):

What is the question

OpenStudy (anonymous):

The equation above

OpenStudy (anonymous):

okay this what i think it is

OpenStudy (anonymous):

3*4^(x/2) = 96 4^(x/2) = 32 Recall that a^(bc) = (a^b)^c. Applying this with 4^(x/2), we get: 4^(x/2) = 32 [4^(1/2)]^x = 32 2^x = 32 Then since 32 = 2^5, we get: 2^x = 32 2^x = 2^5 x = 5 I hope this helps!